OpenStudy (anonymous):

[4.2-2x] - 6.1 if X=0.8

OpenStudy (anonymous):

-3.5

OpenStudy (anonymous):

@usccarolinagurl GIVE the man fish you feed for a day TEACH the man to fish you feed him for a lifetime

OpenStudy (anonymous):

Fine! [4.2-2(0.8)]-6.1=-3.5 0.8x2=1.6 4.2-1.6=2.6 2.6-6.1=-3.5

OpenStudy (anonymous):

TEACH not show him a solved

OpenStudy (anonymous):

haha, nevermind :)